Solve by using substitution.
2x = y - 4
-3y = x - 5

Answer by stanbon(75887)

Solve by using substitution.
2x = y - 4
-3y = x - 5
--------
Rearrange::
2x - y = -4
x +3y = 5
------
Modify for elimination::
2x - y = -4
2x + 6y = 10
----------------
Subtract and solve for "y"::
7y = 14
y = 2
-----
Solve for "x"::
x = -3y+5
x = -6+5 = -1
----------------
Cheers,
Stan H.
-------------